OpenStudy (kitty626oldrose):

WILL FAN AND MEDAL 4 (3x-3y=36) -3(4x-4y=48) if you multiply the first equation by 4 and the second equatian by -3 which variable would be eliminated when you add the equatian together?

OpenStudy (kitty626oldrose):

i got 0=0

OpenStudy (debpriya):

yep I m getting the same

OpenStudy (kitty626oldrose):

12x-12y=144 -12+12y=-144 if i got 0=0 does that mean i have infinitely many solutions

OpenStudy (kitty626oldrose):

@debpriya

OpenStudy (welshfella):

yes infinite solutions

OpenStudy (kitty626oldrose):

@welshfella thank you so muh